Ok so I have 3 matrix' and its A B C and I have to solve (-1 means inverse) for (A-1xB)-1xC so does (A inverse times B) the whole thing inverse times C = (B-1A)C?
1
Expert's answer
2013-02-25T12:15:39-0500
We know that (AB)^(-1) =B^(-1) * A^(-1).
Also, (A^(-1))^(-1)=A.
Now, we can simplify given expression as (A^(-1) * B)^ (-1) * C = (B^(-1) * A)*
C = B^(-1)AC, and this is a result.